OpenStudy (anonymous):

(1+tan)(tan2x)=2tanx/1-tanx I need to prove that the equation is an identity.

OpenStudy (loser66):

tan (2x) = tan (x+x) = \(\dfrac{2tanx}{1-tan^2x}\) replace to tan 2x from the left hand side (1+tan x ) * \(\dfrac{2tanx}{1-tan^2x}\)= right handside